OpenStudy (anonymous):

[4.06]What is the equation of the line that passes through the points (−1, 3) and (1, 11)? 4x − y = 1 4x + y = 1 x − 4y = −13 4x − y = −7

OpenStudy (anonymous):

@Algebraic! @robtobey @SheldonEinstein @Solmyr @dmezzullo @Faris_Waleed @mathteacher1729

Parth (parthkohli):

Point-slope form is classic.

Parth (parthkohli):

\[\rm y-y_1 = m(x - x_1)\]

OpenStudy (anonymous):

I found for the slope 4

OpenStudy (anonymous):

is that correct?

Parth (parthkohli):

Yes!

OpenStudy (anonymous):

the point slope ofrm is y-11=4x(x-1)

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

sorry 4(x-1)

Parth (parthkohli):

\[\rm y - 11 = 4(x - 1)\]Yes!

OpenStudy (anonymous):

yap correct

OpenStudy (anonymous):

then when i solve it gives me y=4x+7

OpenStudy (anonymous):

correct?

Parth (parthkohli):

Right!!

OpenStudy (anonymous):

then i have trouble can you walk me through it?

Parth (parthkohli):

Yes, I will. You have to get all variables on one side.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

show me please?

Parth (parthkohli):

Subtract \(\rm 4x\) from both sides.

OpenStudy (anonymous):

so it gives me -4x-y=7

OpenStudy (anonymous):

?

Parth (parthkohli):

No...\[\rm y = 4x + 7\]\[\rm y - 4x = 4x - 4x + 7\]

OpenStudy (anonymous):

these are my options for the answer 4x − y = 1 4x + y = 1 x − 4y = −13 4x − y = −7

Parth (parthkohli):

The options are wrong. ;)

Parth (parthkohli):

Or, if you need it, try plugging in values.

OpenStudy (anonymous):

no the options are correct I assure you

Parth (parthkohli):

\[\rm 4(1) - 11 = 1 \]the first option is not right.

OpenStudy (anonymous):

may you just give me the answer?

Parth (parthkohli):

\[\rm 4(1) + 11 = 1\]not right either

Parth (parthkohli):

\[1 - 4(11) = -13\]no

OpenStudy (anonymous):

is it the 4 one

Parth (parthkohli):

\[4 - 11 = -7\]

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@ParthKohli can you help me with 2 others?

Parth (parthkohli):

Sure, please post them in a new question... my computer is lagging.

OpenStudy (anonymous):

okay thanks